OpenStudy (jennyrlz):

8(1/3)(cos (216/3) + i sin (216/3)) this turns to 2(cos 72 + i sin 72) how do i turn the 8(1/3) into two?

OpenStudy (anonymous):

DIVIDE Medal?

OpenStudy (jennyrlz):

when i devided i got 2.66(repeating)

OpenStudy (anonymous):

hmmmmmm

OpenStudy (anonymous):

Give me a Min

OpenStudy (jennyrlz):

:D

OpenStudy (anonymous):

Try X And + And -

OpenStudy (anonymous):

Some of those gotta work xD

OpenStudy (jennyrlz):

0.o?

OpenStudy (jennyrlz):

lol xD

OpenStudy (anonymous):

Wait Hold Up

OpenStudy (jennyrlz):

like 216/3 is 72 i understand that

OpenStudy (anonymous):

Let me ask my Superviser

OpenStudy (jennyrlz):

but 8/3 is not 2 >.<

OpenStudy (jennyrlz):

alright

OpenStudy (anonymous):

YEs it is 2.6

OpenStudy (jennyrlz):

but for the answer my teacher got 2 not 2.6 ; ;

OpenStudy (anonymous):

She is Messed up then

OpenStudy (anonymous):

It is 2.6 Not 2 100%

OpenStudy (anonymous):

Im Sorry but there is no other way to see it...

OpenStudy (jennyrlz):

Find the cube roots of 8(cos 216° + i sin 216°). this is the problem

OpenStudy (jennyrlz):

Find the cube roots of 8(cos 216 + i sin 216). the blocks with question marks are replacing the degree symbol

OpenStudy (anonymous):

It's \(8^{1/3}\), not \(8\times\dfrac{1}{3}\).

OpenStudy (jennyrlz):

ohhhhhhhhhhh

OpenStudy (anonymous):

\[8^{1/3}=\sqrt[3]{8}=\sqrt[3]{2^3}=2\]

OpenStudy (jennyrlz):

:DDDD i love you

OpenStudy (jennyrlz):

thanks!! ^-^

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

BUT

OpenStudy (anonymous):

I am in a Company That help students in OpenStudy...